Reports of 61068
AdV-PAY (NE and WE Payloads)
melo - 15:02 Thursday 15 May 2025 (66795) Print this report
Comment to WE thermocamera imaging after recovery (66776)

Here are more pictures of the thermocamera taken today.

Fig. 1: ITF unlocked;

Fig. 2: ITF locked without offset;

Fig. 3: ITF locked with offset applied to WE_Y (12 mm).

No clear difference can be observed between the spots of the two images with the ITF locked. The analysis of the position remains the same of the previous log-entry (#66776).

Images attached to this comment
Detector Operation (Operations Report)
gherardini - 14:59 Thursday 15 May 2025 (66786) Print this report
Operator Report - Morning shift
This morning I found the ITF trying to lock at LOW_NOISE_3, the activity of the morning was focused on west end magnetic injection, magnetic coupling and 3Hz comb investigation; in the afternoon Suzanne started to work on WE and NE mirror thermal camera imagines at CARM_NULL state; the ITF locked/unlocked following the activity, the lock acquisition worked quite well during the shift with some fail during the path to CARM null.
 
others activity done:

- DAQ: investigation on ENV missing channels (#66793);
- CAL: NE newtonian calibrator switch on (#66781);
Images attached to this report
AdV-COM (1/√f noise)
mantovani - 14:45 Thursday 15 May 2025 (66794) Print this report
Comment to Broadband noise after WE mirror cleaning (66790)

Only to have the comparison between the OG in the  various configurations:

- 'Old WE'; 'new WE before cleaning';'new WE before cleaning Y offset';'new WE after cleaning'; 'new WE after cleaning Y offset'

 

Images attached to this comment
AdV-DAQ (Data collection)
masserot - 12:37 Thursday 15 May 2025 (66793) Print this report
DMS grey sometime due to missing ENV channels

The operators reported that from time to time some parts of the DMS became grey due to missing ENV channels  . Several investiagtions were performed

  • By looking at the same SMS channels using dataDisplay, FbmFE level or directly on the FbsMoni server,  ENV SER builder, it appears that the channels are missing only after the FbmFE level and are correctly collected by the FbsMoni server
  • The channels collected by the EnvServer1500W and EnvServerSqz were stored in dedicated SER channels . In these condiftions, there channels were not anymore missing, meaning the issue is related to the ENV SER collection
  • The problems appeared the 2025/05/06 around 7H00 :
    • At the same time period, the SMS data collection of the ENVnoise server was set up for debugging purposes and its SMS channels collected by the FbsISC slow frame builder in a SER called "ENV" too . It was the mistake as a SER is uniq in all the DAQ area , and it 's currenctly built  by the FbsMoni server . The ENVServer SMS data collection from FbsISC has been disabled at  2025-05-15-08h12m54-UTC
    • To avoid this statement, the ENVnoise SMS channels are still collected by the FbsISC server but stored in the ENVnoise and the ENV_NOISE SER channels : operation performed at 2025-05-15-09h10m28-UTC

This should fix the issue

Images attached to this report
AdV-COM (1/√f noise)
sorrentino - 11:39 Thursday 15 May 2025 (66790) Print this report
Broadband noise after WE mirror cleaning

Comparing the sensitivity before and after WE mirror cleaning, see attachment with different curves:

a) red before cleaning, no shift on WE mirror

b) blue before cleaning, shift on WE mirror: optical gain is about 20% better than in a)

c) purple after cleaning, no shift on WE mirror: optical gain is about 5÷6% better than in a)

d) green after cleaning, shift on WE mirror: optical gain is similar as in b)

A quick conclusion is that:

1) the broadband ~1/f^2 noise between 50 and 70 Hz, that was present before cleaning and did not depend on optical gain, is not present anymore

2) the 3 Hz comb, that was present before cleaning and whose ampltude did not depend on optical gain, is not present anymore

3) the broadband (1/f^2/3?) noise between 100 Hz and 200 Hz is scaling with the optical gain; with similar optical gain, it's roughly the same before and after cleaning.

Images attached to this report
Comments to this report:
mantovani - 14:45 Thursday 15 May 2025 (66794) Print this report

Only to have the comparison between the OG in the  various configurations:

- 'Old WE'; 'new WE before cleaning';'new WE before cleaning Y offset';'new WE after cleaning'; 'new WE after cleaning Y offset'

 

Images attached to this comment
AdV-TCS (Ring Heater)
ruggi - 11:22 Thursday 15 May 2025 (66792) Print this report
Comment to WE RH Shutdown and Cable Disconnection for 3 Hz Comb Noise Investigation (66702)

The 50 Hz to be removed from DARM via feed-forward is back to normal. The old startegy for FF parameters computation works again.

Images attached to this comment
AdV-ISC (Automatic Alignment)
mpinto, mantovani - 10:17 Thursday 15 May 2025 (66791) Print this report
ASC noise budget may 15th

we post the angular noise budget computed after a set of measurements performed this morning.

working conditions of the ITF: decentered position of the beam on the WE (12mm offset on vertical).

DIFFp TX (blue curve) shows higher coupling.

Images attached to this report
AdV-SGD (FDS commissioning)
Zendri, De Rossi, Laudenzi, Lunghini - 10:02 Thursday 15 May 2025 (66789) Print this report
Hardware installation for the in fiber phase noise cancellation loop

This is a report on the hardware installations, performed on May 13th,  for  the  phase noise cancellation loop  of the fiber   between Inj and EQB1. The installations involved several laboratories

Atrium:

The fiber circuit of the phase camera and the squeezing in atrium has been modified  (figures 1,2)  in order to decouple the phase camera outputs from the one used to power the box “phase noise box” (fig3) containing the setup for the fiber phase noise cancellation loop . In this way, for the phase noise cancellation loop, we act on a dedicated AOM (operating at 100 MHz) independent from that of the phase camera.

Injection E-Lab:

The electronics for the phase noise cancellation have been installed in the Injection E-room. The installation includes a HP8656 RF generator equipped with a custom RF amplifier that serves to power the AOM at 100 MHz and a new 4-channel DDS generator that provide our low phase noise frequency reference (fig4)

DetLab (EQB1):

The optical fiber coming from the atrium to EQB1 has been split in two. One output goes to the main PLL of the squeezer as in the old set-up and the other branch to the fiber reflector (see fig 1 scheme).

Finally, 3 low attenuation RF cables (L=15 meters) were laid between the atrium and the E-room.

There was no time to test the loop operation, which will be done as soon as possible in agreement with the ITF planning. The feedback system is therefore left off.

Images attached to this report
Non-image files attached to this report
AdV-DAQ (Calibration)
mours - 7:57 Thursday 15 May 2025 (66787) Print this report
Comment to NCal activities on North and West end Buildings (66781)

The NE and NS NCal power supplies were left OFF at the end of the NCal work of yesterday. They have been turned back ON and the NE NCals restarted this morning at 5:42 UTC  

AdV-DAQ (Calibration)
vanhove - 23:28 Wednesday 14 May 2025 (66785) Print this report
Comment to NCal activities on North and West end Buildings (66781)

I updated virgoData/VirgoOnline/SNEB_dbox_rack.cfg to set the reference positions of NW rotors to 0. Attached image shows the position histograms after this operation.

Images attached to this comment
Detector Operation (Operations Report)
amagazzu - 23:09 Wednesday 14 May 2025 (66775) Print this report
Operator Report - Afternoon shift

ITF found in LOCKING_ARMS_BEAT_DRMI_1F, we were able to reach ACQUIRE_LOW_NOISE_3, but the ITF unlocked 13:41 UTC.
From 14:00 UTC to 17:35 UTC the Cal Team worked at the both terminals for the planned NCal Activity (see reports #66781), after that the recovery of the ITF, carried out by Mantovani and Bersanetti, started (see reports #66782, #66784).
Activity concluded at 21:35 UTC, ITF left with AUTORELOCK_FAILSAFE engaged at CARM_NULL_1F.

Images attached to this report
AdV-COM (AdV commissioning (1st part) )
bersanetti, mantovani - 22:42 Wednesday 14 May 2025 (66784) Print this report
Operations on the interferometer

We spent the evening in trying to recover the interferometer, activity which, by the way, we could not do for most of the day.

Despite scheduled times for the different activities, two locks were killed at the end of the acquisition by two ill-advised actions which both impacted SNEB_Fb:

2025-05-14-18h10m23-UTC>ERROR..-[TfbSource::StoreData] producer SNEB_Quadrants_DAQ_10000Hz: overflow at GPS 1431281441.199935050 (received 2001 / expected 2000) (2 times)
2025-05-14-18h10m23-UTC>ERROR..-[TfbSource::StoreData] producer SNEB_LC_DAQ_10000Hz: overflow at GPS 1431281441.199935050 (received 2001 / expected 2000) (47 times)
2025-05-14-18h10m23-UTC>ERROR..-[TfbSource::StoreData] producer SNEB_PSD_sensing_DAQ_10000Hz: overflow at GPS 1431281441.199935050 (received 2001 / expected 2000) (15 times)
2025-05-14-18h10m23-UTC>ERROR..-[TfbSource::StoreData] producer SNEB_Photodiodes_DAQ_10000Hz: overflow at GPS 1431281441.199935050 (received 2001 / expected 2000) (51 times)


2025-05-14-20h27m37-UTC>ERROR..-[TfbSource::StoreData] producer SNEB_Quadrants_DAQ_10000Hz: overflow at GPS 1431289675.199935070 (received 2001 / expected 2000) (2 times)
2025-05-14-20h27m37-UTC>ERROR..-[TfbSource::StoreData] producer SNEB_LC_DAQ_10000Hz: overflow at GPS 1431289675.199935070 (received 2001 / expected 2000) (47 times)
2025-05-14-20h27m37-UTC>ERROR..-[TfbSource::StoreData] producer SNEB_PSD_sensing_DAQ_10000Hz: overflow at GPS 1431289675.199935070 (received 2001 / expected 2000) (15 times)
2025-05-14-20h27m37-UTC>ERROR..-[TfbSource::StoreData] producer SNEB_Photodiodes_DAQ_10000Hz: overflow at GPS 1431289675.199935070 (received 2001 / expected 2000) (51 times)

This sums up to almost two hours lost, of the only five we could work today for recovery.

It is not acceptable not being able to coordinate the work in a decent way. During a lock recovery no other actions should be done on the interferometer.

AdV-ISC (Commissioning up to first full interferometer lock)
mantovani, bersanetti - 22:41 Wednesday 14 May 2025 (66782) Print this report
Lock recovery

We have recovered the LN3 configuration. The recovery went smoothly, reached LN3 at the second attempt after one unlock in locking LN2.

We tested first the configuration in which the beam is centered in the West end mirror.

In figure 1, the comparison between the "after-cleaning" in purple and the "pre-cleaning" in blue sensitivities is shown (beam centered).

- the 3Hz is not present and the low frequency sensitivity is back to the best level

- the frequency noise at high frequency is much higher than before.

We left the ITF in LN3 for ~15 mins and we decided to unlock to test the decentered position of the beam on the WE (12mm offset).

We unlocked at locking in LN3, see next entry.

We leave in auto-relock with the decentering offset on.

Images attached to this report
AdV-DAQ (Calibration)
aubin, dangelser, goeltzenlichter, mours, van-hove - 22:07 Wednesday 14 May 2025 (66783) Print this report
Comment to NCal: installation of West-End setup (2nd day) (66257)

Erratum : the explained axis convention was wrong. We checked it on may 14 2025, here is the correct one:
-Axial Lateral, along the x-axis, corresponding to the laser beam (increasing when moving away from the central building)
-Vertical along the z-axis (positive upward)
-Lateral Axial, along the y-axis, with y = z ^ x

AdV-DAQ (Calibration)
aubin, dangelser, goeltzenlichter, mours, van-hove - 22:05 Wednesday 14 May 2025 (66781) Print this report
NCal activities on North and West end Buildings

Erratum about logbook entry #66257 (https://logbook.virgo-gw.eu/virgo/?r=66257): the explained axis convention was wrong. Here is the correct one:
-Axial Lateral, along the x-axis, corresponding to the laser beam (increasing when moving away from the central building)
-Vertical along the z-axis (positive upward)
-Lateral Axial, along the y-axis, with y = z ^ x
All setups have been checked for conformity with this convention.
All reported times are in UTC.

Work on NCAL in North End Building
13/05/2025
Unload box and start the installation of the new NCal setup around the NE Mirror:
~6:30 Stop all rotors and switch off power supply
-7:13 -9:00 drill hole for new East table
-9:05-10:00 start drilling holes for new West table
-11:10-12:00 finish drilling holes for new West table
- East table screwed on tower base at 18:00
-18:00 Support #2 fixed on East table (NE)
-18:13 zero setting of the NE support

In parallel
- Disassemble “old” NEN and NEF setups:

  • Old NEN (rotor 11 in box5) dismounted. Box was difficult to open so that the covers of box 5 and the axis of rotor 11 need to be “repaired”
  • Old NEF (rotor 3 in box 3) dismounted.
  • All electronic removed (for use in new setup)
  • Support removed.
  • The table remains in place

-Rotor assembly:

  • PVC rotor 16 mounted in box 01 (reusing the stator from rotor 11 box 5, NEN)
  • PVC rotor 17 mounted in box 06 (reusing the stator from rotor 3 box 3, NEF)

Warning : a positioning reference hole of platform #3 are not correct. It must be reamed.

14/05/2025
We continued the NCal activities around the NE mirror:
- we fixed the West table on the floor on the morning
- we put the NW NCal support at it nominal position with reference spacers on the afternoon
- we update the SNEB_dbox_rack configuration to set their reference positions to 0:

  • NE (morning) at gps 1431248200
  • NW (afternoon) at gps 1431278200

- we put the NEN and NWN NCals in place:

  • NEN (morning): box 01 and rotor 16
  • NWN (afternoon): box 06 and rotor 17



Work on NCAL in West End Building
13/05/2025
Cf logbook entry https://logbook.virgo-gw.eu/virgo/?r=66763

14/05/2025

On WSN support, we installed counterweight and rotor 31 (box 11).
On WNN support, we installed counterweight and rotor 33 (box 2).
Electronic cabling done
Setup positions adjusted (see images)

Installation of all West end building rotors is complete. All have been runned between about 16:00 and 17:00 at around 18Hz.

For North End building (NEN and NWN), cabling and alignment must still be completed

Images attached to this report
Comments to this report:
vanhove - 23:28 Wednesday 14 May 2025 (66785) Print this report

I updated virgoData/VirgoOnline/SNEB_dbox_rack.cfg to set the reference positions of NW rotors to 0. Attached image shows the position histograms after this operation.

Images attached to this comment
mours - 7:57 Thursday 15 May 2025 (66787) Print this report

The NE and NS NCal power supplies were left OFF at the end of the NCal work of yesterday. They have been turned back ON and the NE NCals restarted this morning at 5:42 UTC  

AdV-SGD (FDS commissioning)
Martinez, Bouvier, Bonnand, De Laurentis (delauren) - 19:19 Wednesday 14 May 2025 (66779) Print this report
SQB1 FIs temperature control implementtion and setting PART II (May, 14th )

May, 14th, SQB1 FIs -I Ipart

Martinez, Bouvier, Bonnand, De Laurentis

Goal: check the SQB1 FIs temperature control loop.

Befor eto go on with the goal, the yesterday activities has been ended: the connector has been unsoldered to have the right connection (see picture below).

AD_4nXdW5kjR7nRDere3z1J8Jgo4bV9XB7wBdHfo2mzWMkfmK_SIEyVGt6dawWRu3icwEE-MN6jnmvtMWGMCAtXIy9hkp57etd0l8QWMJU6NzjwCSfh4yAX7S_ifA4U78zpZHEZ30EkK?key=RVPlFdqe5si02nNw-d9mZw

Corrected connection plugged to the Peltier controller 

(gray wire : PT100, red wire : Peltier module)

 

It appeared, after testing using a multimeter, that the PT100 of the Faraday 1 is not working. It could be a false contact or a short circuit.

For this reason the control loop test has been made on ly on FI2

ANNOTATION on the cable and DAQ channel

 From labels on the SQB1 HWPs cables in E-room, and from OpticLab in VPM (https://vpmrd.virgo.infn.it:40000/main.html?subsystem=Optic+Lab&process=SQZ_Rot):

  • A corresponds to Axis 1 ->    HWP 1  (toward ITF)

  • B corresponds to Axis 2 →  HWP 2   (betwen FI2 and FI3)

According with this  FI1 and FI2 signals on dataDispaly icorrespond to the cable label A and B, repsectively (see picture below)

Temperature control loop for FI2 test

The control  loop implemented by Alain in the configuration file,  has been  tested using - to start -  the same parameters of the similar OMC temperature control. In practice we copy in the filter file  

lines from 45 to 60 of the OMC temperature control configuration file.
The test was positive but some parameters and configurations (boost seems not working) must be checked. We plan to do this in the next days.

 

 

Images attached to this report
AdV-SGD (FDS commissioning)
Martinez, Bouvier, Bonnand, De Laurentis (delauren) - 19:05 Wednesday 14 May 2025 (66778) Print this report
SQB1 FIs temperature control implementtion and setting PART I (May, 13th )

Goal: check the SQB1 FIs temperature control driver for the Peltier actuators  and the pt100 sensors

  1. Check the driver, yet mounted in the e-room rack (on the top of TCS rack (https://logbook.virgo-gw.eu/virgo/?r=66440https://logbook.virgo-gw.eu/virgo/?r=66440) stand alone with a Peltier module and pt100 sensor and a oscilloscope;

  2. Connect to  the DAC and check the dataDIsplay channels ()https://logbook.virgo-gw.eu/virgo/?r=66598;

  3. Connect to the FI1 and FI2 peltiers and pt100 (cable ready)

First two steps, successufully resched. The last gave a current overload. Checking the possible causes, some swap of the wires on the  connector pins  has been found (see attached picture that compares the pin from the stand alone sensor and actuator and the ones of the connector plugged on the cable coming from SQB1).  In th efollowing picture is it is possible to see:  on the left handside,  the connector plugged on the cable coming from SQB1; on the right handside : the connector of the stand alone sensor and actuator

AD_4nXeomtzdeu7t1xJ9MhwCGARgLRRGWVnM7Ef9rpeqxH-7egtKHaAcG3GG9PYnv5gHgcPiR78V8uib6bpQGUwsgKKWZcH7PEQVt_j8SXllMZ8DFkM-zKX2e6qJW5_-4Ih-FY1xlP6a?key=RVPlFdqe5si02nNw-d9mZw

 

AdV-TCS (Ring Heater)
fiori, tringali, paoletti - 17:02 Wednesday 14 May 2025 (66777) Print this report
Comment to WE RH Shutdown and Cable Disconnection for 3 Hz Comb Noise Investigation (66702)

In today's lock the 3Hz comb and the broadband noise underneath seem no more present. Also the 178Hz peak is (at least reduced).

Images attached to this comment
AdV-PAY (NE and WE Payloads)
melo, cagnoli - 16:47 Wednesday 14 May 2025 (66776) Print this report
WE thermocamera imaging after recovery

While the ITF was locking this afternoon I took some images with the thermocamera installed on WE viewport.

Fig. 1: shows the ITF unlocked.

Fig. 2: shows the ITF locked in CARM_NULL.

Fig. 3: shows the reconstructed image including the baffles done by G. Cagnoli (red one is the inner baffle, cyan one is the outter baffle, the green circle is the mirror surface with 2cm grid lines and the orange sign is the center of the mirror) superposed to the thermocamera images. I also put a red dot to highlight the bright spot seen by thermocamera. 

To be kept in mind for comparison when we will be able to do more tests with different offsets.

Images attached to this report
Comments to this report:
melo - 15:02 Thursday 15 May 2025 (66795) Print this report

Here are more pictures of the thermocamera taken today.

Fig. 1: ITF unlocked;

Fig. 2: ITF locked without offset;

Fig. 3: ITF locked with offset applied to WE_Y (12 mm).

No clear difference can be observed between the spots of the two images with the ITF locked. The analysis of the position remains the same of the previous log-entry (#66776).

Images attached to this comment
AdV-ISC (Commissioning up to first full interferometer lock)
boldrini, spinicelli, bersanetti, menzione - 16:24 Wednesday 14 May 2025 (66772) Print this report
lock acquisition recovery

After Boschi recovered the WE inertial damping, we proceeded with the recovery.

The ITF reached LN2 by itself, but suffered un unlock, most likely due to a drift on BS_TX. Upon further investigation, we discovered that the galvo loop for B4_QD2, used for the high frequency branch of the error signal, was not engaged (Fig.1). This specific galvo loop is never engaged by Metatron, it probably never posed any problem. We manually re-engaged the loop and added a command in ITF_LOCK.py, line 3909, to engage it before switching the BS_TX error signal.

The next attempt was disrupted by an unlock in CARM_NULL_1F, which left the WI guardian open. We investigated and discovered that the suspension's top stage tilted by ~800 urad on TY in about half a second before the unlock (Fig.2). Ruggi and Pinto recovered the WI suspension and are investigating the issue. The top stage control disengaged due to a corrupted data package sent to the dsp.

The third attempt successfully reached ACQUIRE_LN3, but the ITF unlocked after a glitch on NE LN correction (Fig.3). Even though the acquisition of LN3 was not complete, the DCP was just ~20 Hz away from the threshold that triggers the transition and the sensitivity was about 32 Mpc.

In agreement with Mantovani and Was, we allow the work of Ncal for the moment.

Images attached to this report
Detector Operation (Operations Report)
menzione - 14:56 Wednesday 14 May 2025 (66767) Print this report
Operator Report - Morning shift

ITF found in LOCKED_ARMS_IR in UPGRADING mode.

- WE ID loop recovery concluded (Ruggi, Pinto)
- Lock recovery in progress...
- SQB1 Faraday temperature controller (De Laurentis) in progress...
NE Ncal support (Mours) in progress... In pause up to 14:00 UTC

At 11:00 UTC I started the lock acquisition (COMMISSIONING mode set).
After the usual cross-alignment in ACQUIRE_DRMI I locked ITF up to LN2 at first attempt. During the lock I re-enabled again SDB2 floating setpoint.

Images attached to this report
AdV-DAQ (Calibration)
aubin, dangelser, goetzenlichter, mours, van-hove - 14:05 Wednesday 14 May 2025 (66758) Print this report
NCal activities: Unbalance rotor 32 WEB (WEN)
First NCal modification12/05/2025 afternoon at 18:00 utc:

We change the unbalance correction of rotor 32 (WEB: WEN) on the "bearing side" (opposite to motor side) by an half-ring screwed on holes 2 and 3 (IMG_8114)
Previous unbalance had its cutted sided on the side of the screw #4 (see IMG_8108)
Images attached to this report
AdV-SAT (Suspension control upgrade)
pinto, ruggi - 12:55 Wednesday 14 May 2025 (66770) Print this report
Comment to WE ACCH1 re-calibrated (66595)

Today we investigated the issue experienced yesterday with the ID of the WE.

We performed a measument in order to recompute the calibration for the accelerometeres. We found out that ACC_H1 was the one miscalibrated, but the correct compensator was the one used before the last update. One possibility is that the accelerometer switches between two states. The reason of this behaviour is not clear. Some investigation is needed.

In the meantime, we reloaded the older compensation filter, and we modified the DSP card adding one switch for the two compensators in order to have the possibility to move between one 'state' and the other.

AdV-ISC (Commissioning up to first full interferometer lock)
mwas - 8:34 Wednesday 14 May 2025 (66768) Print this report
Comment to ITF recovery (66765)

This is a good reminder.

 Figure 1. shows that SDB1 has changed orientation compared to where it was a few weeks ago, and during the recovery it has moved further away.

This morning I have disabled the floating set points (the bench learning and remembering on its own what are the set points), and put back the set points to the average values of the bench orientation in TX and TY from beginning of May.

We will see if helps, the interferometer locked on DRMI for 20 seconds, and the B1p quadrants galvo close during that time.

Images attached to this comment
Detector Operation (Operations Report)
gherardini - 22:55 Tuesday 13 May 2025 (66762) Print this report
Operator Report - Afternoon shift

The work on newtonian calibrators, both on north end and west end, went on in the afternoon in parallel with the west arm lock recovery; the west arm has been realigned and relocked with the green lock too, during this work we noted a strange behavior of the WE suspension control with a slow oscillation that grow up, switching off the ID control it disappeared; I called Valerio to work on this problem, he didn't able to solve it postponing any other intervention to tomorrow morning, we continued to work with the WE controlled without the inertial damping; the work on newtonian calibrator stopped at around 18:15UTC and we continued with the lock recovery, we succeded to lock up to the DRMI but the lock was not good; we stopped here, I left the cavities locked on the infrared. Oncall events

Oncall events

SUSP
(13-05-2025 15:15 - 13-05-2025 16:00) From remote
Status: Ended
Description: WE inertial damping not working

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×